# Fermi-Energy-Sensitive Universal Conductance Fluctuations in Anisotropic Materials

### v2

@@ I Introduction
Universal conductance fluctuation (UCF) manifests itself as aperiodic and reproducible fluctuations superimposed on the ohmic resistance of disordered mesoscopic devices when scanning the gate voltage or external magnetic field [1, 2, 3]. The aperiodic fluctuations in diffusive transport stem from the fundamental quantum interference among all the phase-coherent Feymann paths of the traversing electrons. Thus, UCF serves as fingerprints for metallic materials due to their sensitivity to complex impurity configurations and sample-specific characteristics [4, 5, 6]. Despite that, the Altshuler-Lee-Stone theory has established a connection between the universal statistical properties of conductance fluctuations and the valuable information of materials, such as the dimensions and symmetry classes [7, 8, 9]. For example, in a standard two-terminal setup, diagrammatic calculations using a free electron-gas model have found that the intrinsic amplitude of UCF at zero-temperature is determined by the number of uncorrelated bands $k$, the level degeneracy $s$, the symmetry parameter $\beta$, as well as the shape of the material, which reads [7]:
@@
in units of $\frac{e^{2}}{h}$. Here $\beta=1,2,4$ for orthogonal, unitary and symplectic systems, respectively, and the effect of dimension is incorporated in the prefactor $c_{d}$. Equation  1 highlights the UCF in characterizing the materials’ symmetry class as well as their dimension properties in the diffusive regime. Its value is universal and independent of the materials’ details, such as the magnitude of conductance, system size, Fermi energy, disorder strength, and the strength of magnetic fields applied to samples.
According to Eq. 1, the integers $k,s,\beta$ and the prefactor $c_{d}$ contribute to the UCF amplitude independently. The formers’ contribution is evidenced by the experimentally observed decay of the UCF amplitude upon the application of an external magnetic field, which suppresses the Cooperon mode and turns $\beta=1$ or $4$ into $\beta=2$ [10, 11, 12, 13, 14]. The latter contribution from $c_{d}$ has a much more complex parameter dependence. Before touching on the main focus of our work on the Fermi energy dependence of $c_{d}$, we briefly review several known aspects. In an isotropic free electron gas considered by the Altshuler-Lee-Stone theory, the coefficient $c_{d}$ is a function of the system size $L_{\alpha}(\alpha=x,y,x)$ for metallic samples. It becomes quite sensitive to $L_{x/y}$ when the transversal directions $L_{x/y}$ are longer than the longitudinal transport direction $L_{z}$, and diverges in the limit of $L_{x/y}\gg L_{z}$ [15, 16]. Moreover, the value of $c_{d}$ also varies significantly for systems beyond the isotropic free-electron gas model. Further developments of the theory have addressed several aspects of interest in the conductance fluctuations, including the proximity to a superconductor [17, 18, 19], the positions of leads and geometry of the probing devices [20, 21], boundary conditions [16], the non-integer dimensions [22, 23], the near-ballistic regime [24], spin-orbital coupling [25, 26, 27, 28] and band anisotropy [15], etc. Despite the factors that modify $c_{d}$, it is generally considered as an irrelevant constant of order unity for a given material in the literature for several reasons. First, for a given experimental device or material, it is not easy to calculate the specific value of $c_{d}$ using realistic parameters. Instead, applying the Altshuler-Lee-Stone theory to estimate its magnitude in experiments is convenient because of its simplicity and broad applicability. Second, the relative variance of $c_{d}$ is usually not sensitive to parameters such as Fermi energy or magnetic field, partially because these fields only vary within a small range. Third, for most experimentally relevant quasi-1D nanowire devices, $c_{d}$ is pinned at 0.365, which is robust against other perturbations. Above all, the role of $c_{d}$ is not as much emphasized in most literature.
@@
The experimental observation [51] raises a natural and unaddressed question about the sensitivity of $c_{d}$ to Fermi energy in anisotropic materials. This question is relevant to future studies on UCF because anisotropic materials are abundant in nature. Advanced technologies are well developed to engineer anisotropy through strain or twist  [58, 59, 60, 61, 62, 43, 63]. For example, the UCF of the twisted bilayer graphene quantum dot  [63] shows a negligible dependence on the twisting angles, suggesting that materials with stronger anisotropy than graphene are necessary for a UCF dependent on the twist angle. This is also a nontrivial problem because it contradicts popular beliefs that UCF is Fermi-energy independent in general, according to the Althsuler-Lee-Stone theory.
@@
Starting from a continuous free electron-gas model, we analytically determine that $c_{d}$ in Eq. 1 should be replaced by $\tilde{c}_{d}(t_{x},t_{y},t_{z})$. This new term is a function of the diffusion time $t_{\alpha}$ $(\alpha=x,y,x)$ along the dimension $L_{\alpha}$ in the presence of anisotropic dispersion. Numerical evidence and physical arguments are provided for the analytic results. Our analysis reveals an equivalence between the band anisotropy and the spatial anisotropy in $L_{\alpha}$. Identifies regions of the anisotropy parameter where $\tilde{c}_{d}$ is sensitive or insensitive to the band anisotropy. Then, we show numerically that, in the parameter regions of sensitivity, the coefficient $\tilde{c}_{d}$, and thus the UCF amplitude is sensitive to the Fermi energy, which doubles the UCF for realistic anisotropy parameters accessible to ordinary materials. These conclusions hold for 2D and 3D materials. Finally, we use realistic material parameters for $\mathrm{Cd_{3}As_{2}}$, and demonstrate the Fermi energy dependence of UCF, in consistent with the experiments. We also find that the increase of the UCF versus Fermi energy is a robust feature in $\mathrm{Cd_{3}As_{2}}$, not necessarily requiring transport along the $L_{z}$ direction, which was assumed in the previous theory [51, 53]. These results should be amenable to future experimental verifications. Combined with the well-developed method to tune the Fermi energy through gate voltages, our findings will allow for delicate control of the UCF in future device design using materials with anisotropic band structures.
@@ II Model and method
Next, we discretize the continuum model into a lattice system as follows [64]
@@
To study the electronic transport, we consider a two-terminal device consisted with a central region and two semi-infinitely long leads. The leads can be modeled by $H_{0}$, whereas the central region is modeled by: $H^{cen}=H_{0}+U(\mathbf{r})$, where $U(\mathbf{r})=\sum_{\mathbf{r}}\varepsilon(\mathbf{r})a_{\mathbf{r}}^{\dagger}a_{\mathbf{r}}$ is the uniform on-site random potentials, with $\varepsilon(\mathbf{r})$ distributed uniformly within $[\frac{-W}{2},\frac{W}{2}]$, and $W$ the disorder strength. Throughout this paper, we fix $M_{z}=0.5$ and $a=b=c=1$, with the unit of energy and length set to unity unless otherwise noted. For simplicity, we neglect the spin degree of freedom. The band edges of Eq. 3 are zero energy at the bottom and $4(M_{x}+M_{y}+M_{z})$ at the top, respectively. A discussion of its Fermi surface and density of states is presented in Appendix  B.
Numerically, we compute the zero-temperature conductance by the non-equilibrium Green’s function method [64, 65, 66, 67, 68, 69, 70] $G=\frac{e^{2}}{h}{\rm Tr}[\Gamma_{L}G^{r}\Gamma_{R}G^{a}]$, where $G^{r/a}(E_{F})=[(E_{F}\pm i\eta){\textbf{I}}-{\textbf{H}}^{cen}-\Sigma_{L}-\Sigma_{R}]^{-1}$ is the retarded(advanced) Green function. ${\textbf{H}}^{cen}$ is the Hamiltonian matrix of the central region, I is the identity matrix, $\eta$ is an infinitesimal positive number, $\Gamma_{L/R}=i(\Sigma_{L/R}^{r}-\Sigma_{L/R}^{a})$ is the linewidth function, and $\Sigma_{L/R}$ is the self-energy of the left (right) lead. In our work, ${\textbf{H}}^{cen}$ is taken to be either Eq. 3 or Eq. 13 below with disorder. The conductance fluctuation $\Delta G$ is the standard deviation of conductance $G$, defined as $\Delta G=\sqrt{\langle(G-\langle G\rangle)^{2}\rangle}$, calculated over an ensemble of more than 1200 disorder configurations.
@@ III.1 Theoretical results of UCF for anisotropic band dispersion
Figure 2: (Color online) The shapes of the Fermi surface in the momentum space $(k_{x},k_{y},k_{z})$, calculated for the (a) continuous Hamiltonian in Eq. 2 and (b-d) discretized Hamiltonian in Eq. 3, as well as their projections onto the $k_{y}-k_{z}$ plane. $M_{z}=0.5$ is fixed, and $M_{x}=M_{y}$. (a-1)-(a-6) The Fermi energy $E_{F}=0.4$ is fixed for different anisotropic parameters $\frac{M_{z}}{M_{y}}=5.0,2.0,1.0,0.5,0.2,0.05$. (b-1)-(b-6) The same parameters as in (a) but for the discretized Hamiltonian. (c-1)-(c-6) The anisotropy parameter $\frac{M_{z}}{M_{y}}=0.2$ is fixed for different reduced Fermi energy $E^{r}_{F}=\Delta,2\Delta,4\Delta,5\Delta,6\Delta$ and $8\Delta$, respectively. $\Delta=(M_{x}+M_{y}+M_{z})/4$ such that the bandwidth is $16\Delta$ and $8\Delta$ is at the band center. (d-1)-(d-6) Same as in (c) but for the anisotropy parameter $\frac{M_{z}}{M_{y}}=5.0$. In (a-d), the labeling ticks of the axis are the same as that in (a-1).
In this section, we obtain the UCF for an anisotropic band by following Ref. [7], whose starting point is the impurity-averaged perturbation theory and does not assume the isotropy of the system as a precondition. Without going into the details of the derivations, we list several key modifications due to anisotropic dispersions and use the existing result in Ref. [7] to simplify the deduction.
@@
where we have used the relation $t_{\alpha}=L_{\alpha}^{2}/(2D_{\alpha})$ in the last line. As is shown in the Appendix  A, the reduced eigenvalues $\tilde{\lambda}_{n}$ in Eq. 8 are sufficient for determining $c_{d}$ in Eq. 1, which in turn gives the UCF amplitude $\Delta G$.
The comments on each line in Eq. 8 are presented as follows. For isotropic systems with $D_{\alpha}=D$, the prefactor $c_{d}(L_{x},L_{y},L_{z})$ is a function of the ratios of the transverse system sizes $L_{x}$ and $L_{y}$ over the longitude system size $L_{z}$ in the transport direction. This is a well-known fact of the spatial dimension dependence of UCF. For anisotropic systems, however, the anisotropic diffusion vector effectively modifies the $L_{\alpha}$ dependence, and $c_{d}$ is replaced by $\tilde{c}_{d}(t_{x},t_{y},t_{z})$ that explicitly depends on the ratios of diffusion time across the three dimensions of the sample.
@@
The theoretical results of $\Delta G=\tilde{c}_{d}(t_{x},t_{y},t_{z})$ versus combinations of $(M_{y},M_{z})$ for $M_{y},M_{z}\in[0,3]$ have been calculated [see Appendix  A for details] and plotted in Fig. 1. It can be seen that when $M_{z}/M_{y}<1$, the UCF amplitude approaches $0.365$, the quai-1D value for $c_{d}$. When $M_{z}/M_{y}>1$, the UCF amplitude increases rapidly and shoots up by more than ten times around the corner $(0,3)$ in the $M_{y}-M_{z}$ plane. This indicates that the UCF amplitude can be extremely sensitive to band anisotropy and will cause significantly measurable effects in experiments.
@@
The above scenario is similar to the sensitivity of $c_{d}$ to the system size $L_{\alpha}(\alpha=x,y,x)$. For transport along z direction, the typical behavior of $c_{d}$ is as follows. For samples of quasi-1D shape, $c_{d}$ is pinned around $0.365$ and is insensitive to the transverse sizes $L_{x,y}$ as long as $L_{z}/L_{x,y}\gg 1$. Its value slightly increases to 0.431 in qusi-2D ($L_{z}/L_{x}=L_{y}/L_{x}\gg 1$) and 0.548 in quasi-3D ($L_{z}=L_{y}=L_{x}$) samples [7, 15]. But $c_{d}$ becomes extremely sensitive to the transverse sizes once the sample is wide enough ($L_{z}/L_{x,y}<1$) and diverges for ($L_{z}/L_{x,y}\ll 1$[15, 16].
@@
To visualize the band anisotropy in momentum space for several typical cases in Fig. 1, we plot the Fermi surfaces of the free electron gas in Eq. 2 in Fig. 2(a-1)-(a-6). The conductance fluctuations continuously decrease from the enhanced value in (a-1) to a minimal quasi-1D value in (a-6), as the shape of the Fermi surface also approaches a quasi-1D ellipsoid in momentum space. For $\frac{M_{z}}{M_{y}}>1$ in Fig. 2(a-1)-(a-2), the Fermi surface is compressed along the transport direction. The diffusion along the $L_{z}$ direction takes a much shorter time $t_{z}$ than $t_{x/y}$ along the transverse direction due to the larger diffusion coefficient $D_{z}>D_{x/y}$. Presumably, one can picture that in the limit of $t_{z}/t_{x/y}\rightarrow 0$, the electrons fly through the sample without fully self-average along the transverse paths, thus leading to conductances with more significant variance. In contrast, when $\frac{M_{z}}{M_{y}}<1$ in Fig. 2(a-4)-(a-6), the Fermi surface is stretched along the transport direction, forming a quasi-1D ellipsoid. The diffusion along the $L_{z}$ direction takes a much longer time $t_{z}$ than $t_{x/y}$ because the electron diffuses slower along the transport direction, i.e., $D_{z}<D_{x/y}$. In the limit of $t_{z}/t_{x/y}\rightarrow\infty$, the electrons have fully traversed the $L_{x/y}$ several times before leaving the central region. Thus, the transmission after ergodic scattering shows only minimal conductance fluctuations.
@@
We conclude this section by noting the equivalent roles of the spatial size $L_{\alpha}$ and the anisotropic parameters $M_{\alpha}\propto D_{\alpha}$ in Eq. 9. Another way to understand the equivalence is to absorb the anisotropy parameters in the Hamiltonian in Eq. 2 into the Hamiltonian in Eq. 3 by rescaling the lattice constant in the hopping term in Eq. 5 as $\hat{a}_{\alpha}\equiv a_{\alpha}/\sqrt{M_{\alpha}}$. This way, one can interpret that the system has isotropic dispersion but with anisotropic hopping terms due to the choice of lattice constants. Meanwhile, the system size under this interpretation is rescaled to $\hat{L}_{\alpha}\equiv L_{\alpha}/\sqrt{M_{\alpha}}$, that is, the same number of lattice sites multiplied by the lattice constant $\hat{a}_{\alpha}$. The fact that band anisotropy and spatial anisotropy are transferrable justifies their equivalence. In analogy with the dependence of $c_{d}$ on spatial anisotropy, we define the regions of sensitivity ($M_{z}/M_{y}\geq 1$) and insensitivity($M_{z}/M_{y}<1$) for $\tilde{c}_{d}$ in Fig. 1, as separated by the dashed line. Within the region of (in)sensitivity, $\tilde{c}_{d}$ is (in)sensitive to the changes in anisotropy parameter $M_{z}/M_{y}$. In summary, in the presence of both anisotropies in $L_{\alpha}$ and $M_{\alpha}$, it is convenient to use the relative diffusion time $t_{z}/t_{x/y}$ that converts all these anisotropies to estimate the UCF amplitude in the diffusive free-electron gas.

### v3

# Fermi Energy Sensitive Universal Conductance Fluctuations in Anisotropic Materials

@@ I Introduction
Universal conductance fluctuation (UCF) manifests itself as aperiodic and reproducible fluctuations superimposed on the ohmic resistance of disordered mesoscopic devices when scanning the gate voltage or external magnetic field [1, 2, 3]. The aperiodic fluctuations in diffusive transport stem from the fundamental quantum interference among all the phase-coherent Feynman paths of the traversing electrons. Thus, UCF serves as fingerprints for metallic materials due to their sensitivity to complex impurity configurations and sample-specific characteristics [4, 5, 6]. Despite that, the Altshuler-Lee-Stone theory has established a connection between the universal statistical properties of conductance fluctuations and the valuable information of materials, such as the dimensions and symmetry classes [7, 8, 9]. For example, in a standard two-terminal setup, diagrammatic calculations using a free electron-gas model have found that the intrinsic amplitude of UCF at zero-temperature is determined by the number of uncorrelated bands $k$, the level degeneracy $s$, the symmetry parameter $\beta$, as well as the shape of the material, which reads [7]:
@@
in units of $\frac{e^{2}}{h}$. Here $\beta=1,2,4$ for orthogonal, unitary and symplectic systems, respectively, and the effect of dimension is incorporated in the prefactor $c_{d}$. Equation  1 highlights the UCF in characterizing the materials’ symmetry class and their dimension properties in the diffusive regime. Its value is universal and independent of the materials’ details, such as the magnitude of conductance, system size, Fermi energy, disorder strength, and the strength of magnetic fields applied to samples.
According to Eq. 1, the integers $k,s,\beta$ and the prefactor $c_{d}$ contribute to the UCF amplitude independently. The formers’ contribution is evidenced by the experimentally observed decay of the UCF amplitude upon the application of an external magnetic field, which suppresses the Cooperon mode and turns $\beta=1$ or $4$ into $\beta=2$ [10, 11, 12, 13, 14]. The latter contribution from $c_{d}$ has a much more complex parameter dependence. Before touching on the main focus of our work on the Fermi energy dependence of $c_{d}$, we briefly review several known aspects. In an isotropic free electron gas considered by the Altshuler-Lee-Stone theory, the coefficient $c_{d}$ is a function of the system size $L_{\alpha}(\alpha=x,y,z)$ for metallic samples. It becomes quite sensitive to $L_{x/y}$ when the transversal directions $L_{x/y}$ are longer than the longitudinal transport direction $L_{z}$, and diverges in the limit of $L_{x/y}\gg L_{z}$ [15, 16]. Moreover, the value of $c_{d}$ also varies significantly for systems beyond the isotropic free-electron gas model. Further developments of the theory have addressed several aspects of interest in the conductance fluctuations, including the proximity to a superconductor [17, 18, 19], the positions of leads and geometry of the probing devices [20, 21], boundary conditions [16], the non-integer dimensions [22, 23], the near-ballistic regime [24], spin-orbital coupling [25, 26, 27, 28] and band anisotropy [15], etc. Despite the factors that modify $c_{d}$, it is generally considered as an irrelevant constant of order unity for a given material in the literature for several reasons. First, for a given experimental device or material, it is not easy to calculate the specific value of $c_{d}$ using realistic parameters. Instead, applying the Altshuler-Lee-Stone theory to estimate its magnitude in experiments is convenient because of its simplicity and broad applicability. Second, the relative variance of $c_{d}$ is usually not sensitive to parameters such as Fermi energy or magnetic field, partially because these fields only vary within a small range. Third, for most experimentally relevant quasi-1D nanowire devices, $c_{d}$ is pinned at 0.365, which is robust against other perturbations. Above all, the role of $c_{d}$ is not as much emphasized in most of the literature.
@@
The experimental observation [51] raises a natural and unaddressed question about the sensitivity of $c_{d}$ to Fermi energy in anisotropic materials. This question is relevant to future studies on UCF because anisotropic materials are abundant in nature. Advanced technologies are well developed to engineer anisotropy through strain or twist  [58, 59, 60, 61, 62, 43, 63]. For example, the UCF of the twisted bilayer graphene quantum dot  [63] shows a negligible dependence on the twisting angles, suggesting that materials with stronger anisotropy than graphene are necessary for a UCF dependent on the twist angle. This is also a nontrivial problem because it contradicts popular beliefs that the UCF is Fermi energy independent in general, according to the Altshuler-Lee-Stone theory.
@@
Starting from a continuous free electron-gas model, we analytically determine that $c_{d}$ in Eq. 1 should be replaced by $\tilde{c}_{d}(t_{x},t_{y},t_{z})$. This new term is a function of the diffusion time $t_{\alpha}$ $(\alpha=x,y,x)$ along the dimension $L_{\alpha}$ in the presence of anisotropic dispersion. Numerical evidence and physical arguments are provided for the analytic results. Our analysis reveals an equivalence between the band anisotropy and the spatial anisotropy in $L_{\alpha}$, and identifies regions of the anisotropy parameter where $\tilde{c}_{d}$ is sensitive or insensitive to the band anisotropy. Then, we show numerically that, in the parameter regions of sensitivity, the coefficient $\tilde{c}_{d}$, and thus the UCF amplitude is sensitive to the Fermi energy, which doubles the UCF for realistic anisotropy parameters accessible to ordinary materials. These conclusions hold for 2D and 3D materials. Finally, we use realistic material parameters for $\mathrm{Cd_{3}As_{2}}$, and demonstrate the Fermi energy dependence of UCF, in consistent with the experiments. We also find that the increase of the UCF versus Fermi energy is a robust feature in $\mathrm{Cd_{3}As_{2}}$, not necessarily requiring transport along the $L_{z}$ direction, which was assumed in the previous theory [51, 53]. These results should be amenable to future experimental verifications. Combined with the well-developed method to tune the Fermi energy through gate voltages, our findings will allow for delicate control of the UCF in future device design using materials with anisotropic band structures.
@@ II Model and method
Next, we discretize the continuum model into a lattice system as follows
@@
To study electronic transport, we consider a two-terminal device consisting of a central region and two semi-infinite leads. The leads can be modeled by $H_{0}$, whereas the central region is modeled by: $H^{cen}=H_{0}+U(\mathbf{r})$, where $U(\mathbf{r})=\sum_{\mathbf{r}}\varepsilon(\mathbf{r})a_{\mathbf{r}}^{\dagger}a_{\mathbf{r}}$ are the uniform random potentials on site, with $\varepsilon(\mathbf{r})$ distributed uniformly within $[\frac{-W}{2},\frac{W}{2}]$, and $W$ the strength of the disorder. Throughout this paper, we fix $M_{z}=0.5$ and $a=b=c=1$, with the unit of energy and length set to unity unless otherwise noted. For simplicity, we neglect the spin degree of freedom. The band edges of Eq. 3 are zero energy at the bottom and $4(M_{x}+M_{y}+M_{z})$ at the top, respectively. A discussion of its Fermi surface and density of states is presented in Appendix  B.
[…]
$G=\frac{e^{2}}{h}{\rm Tr}[\Gamma_{L}G^{r}\Gamma_{R}G^{a}]$ (6)
where $G^{r/a}(E_{F})=[(E_{F}\pm i\eta){\textbf{I}}-{\textbf{H}}^{cen}-\Sigma_{L}-\Sigma_{R}]^{-1}$ is the retarded(advanced) Green function. ${\textbf{H}}^{cen}$ is the Hamiltonian matrix of the central region, I is the identity matrix, $\eta$ is an infinitesimal positive number, $\Gamma_{L/R}=i(\Sigma_{L/R}^{r}-\Sigma_{L/R}^{a})$ is the line width function, and $\Sigma_{L/R}$ is the self-energy of the left (right) lead. In our work, ${\textbf{H}}^{cen}$ is taken to be either Eq. 3 or Eq. 14 below with disorder. The conductance fluctuation $\Delta G$ is the standard deviation of conductance $G$, defined as $\Delta G=\sqrt{\langle(G-\langle G\rangle)^{2}\rangle}$, calculated over an ensemble of more than 1200 disorder configurations.
@@ III.1 Theoretical results of UCF for anisotropic band dispersion
Figure 2: (Color online) The shapes of the Fermi surface in the momentum space $(k_{x},k_{y},k_{z})$, calculated for the (a) continuous Hamiltonian in Eq. 2 and (b-d) discretized Hamiltonian in Eq. 3, as well as their projections onto the $k_{y}-k_{z}$ plane. $M_{z}=0.5$ is fixed, and $M_{x}=M_{y}$. (a-1)-(a-6) The Fermi energy $E_{F}=0.4$ is fixed for different anisotropic parameters $\frac{M_{z}}{M_{y}}=5.0,2.0,1.0,0.5,0.2,0.05$. (b-1)-(b-6) The same parameters as in (a) but for the discretized Hamiltonian. (c-1)-(c-6) The anisotropy parameter $\frac{M_{z}}{M_{y}}=0.2$ is fixed for different reduced Fermi energy $E^{r}_{F}=\Delta,2\Delta,4\Delta,5\Delta,6\Delta$ and $8\Delta$, respectively. $\Delta=(M_{x}+M_{y}+M_{z})/4$ such that the bandwidth is $16\Delta$ and $8\Delta$ is at the band center. (d-1)-(d-6) Same as in (c) but for the anisotropy parameter $\frac{M_{z}}{M_{y}}=5.0$. In (a-d), the labeling ticks of the axis are the same as those in (a-1).
In this section, we obtain the UCF for an anisotropic band following Ref. [7], whose starting point is the impurity-averaged perturbation theory and does not assume the isotropy of the system as a precondition. Without going into the details of the derivations, we list several key modifications due to anisotropic dispersions and use the existing result in Ref. [7] to simplify the deduction.
@@
where we have used the relation $t_{\alpha}=L_{\alpha}^{2}/(2D_{\alpha})$ in the last line. As shown in Appendix  A, the reduced eigenvalues $\tilde{\lambda}_{n}$ in Eq. 9 are sufficient to determine $c_{d}$ in Eq. 1, which in turn gives the UCF amplitude $\Delta G$.
The comments on each line in Eq. 9 are presented as follows. For isotropic systems with $D_{\alpha}=D$, the prefactor $c_{d}(L_{x},L_{y},L_{z})$ is a function of the ratios of the transverse system sizes $L_{x}$ and $L_{y}$ over the longitudinal system size $L_{z}$ in the transport direction. This is a well-known fact of the spatial dimension dependence of UCF. For anisotropic systems, however, the anisotropic diffusion vector effectively modifies the $L_{\alpha}$ dependence, and $c_{d}$ is replaced by $\tilde{c}_{d}(t_{x},t_{y},t_{z})$ that explicitly depends on the ratios of diffusion time across the three dimensions of the sample.
@@
The theoretical results of $\Delta G=\tilde{c}_{d}(t_{x},t_{y},t_{z})$ versus combinations of $(M_{y},M_{z})$ for $M_{y},M_{z}\in[0,3]$ have been calculated [see Appendix  A for details] and plotted in Fig. 1. It can be seen that when $M_{z}/M_{y}<1$, the UCF amplitude approaches $0.365$, the quasi-1D value for $c_{d}$. When $M_{z}/M_{y}>1$, the UCF amplitude increases rapidly and shoots up by more than ten times around the corner $(0,3)$ in the $M_{y}-M_{z}$ plane. This indicates that the UCF amplitude can be extremely sensitive to band anisotropy and will cause significantly measurable effects in experiments.
@@
The above scenario is similar to the sensitivity of $c_{d}$ to the system size $L_{\alpha}(\alpha=x,y,z)$. For transport along z direction, the typical behavior of $c_{d}$ is as follows. For samples of quasi-1D shape, $c_{d}$ is pinned around $0.365$ and is insensitive to the transverse sizes $L_{x,y}$ as long as $L_{z}/L_{x,y}\gg 1$. Its value slightly increases to 0.431 in quasi-2D ($L_{z}/L_{x}=L_{y}/L_{x}\gg 1$) and 0.548 in quasi-3D ($L_{z}=L_{y}=L_{x}$) samples [7, 15]. But $c_{d}$ becomes extremely sensitive to the transverse sizes once the sample is wide enough ($L_{z}/L_{x,y}<1$) and diverges for ($L_{z}/L_{x,y}\ll 1$[15, 16].
@@
To visualize the band anisotropy in momentum space for several typical cases in Fig. 1, we plot the Fermi surfaces of the free electron gas in Eq. 2 in Fig. 2(a-1)-(a-6). The conductance fluctuations continuously decrease from the enhanced value in (a-1) to a minimal quasi-1D value in (a-6), as the shape of the Fermi surface also approaches a quasi-1D ellipsoid in momentum space. For $\frac{M_{z}}{M_{y}}>1$ in Fig. 2(a-1)-(a-2), the Fermi surface is compressed along the transport direction. The diffusion along the $L_{z}$ direction takes much shorter $t_{z}$ than $t_{x/y}$ along the transverse direction due to the larger diffusion coefficient $D_{z}>D_{x/y}$. Presumably, one can picture that in the limit of $t_{z}/t_{x/y}\rightarrow 0$, the electrons fly through the sample without fully self-averaging along the transverse paths, thus leading to conductances with more significant variance. In contrast, when $\frac{M_{z}}{M_{y}}<1$ in Fig. 2(a-4)-(a-6), the Fermi surface is stretched along the transport direction, forming a quasi-1D ellipsoid. The diffusion along the $L_{z}$ direction takes a much longer time $t_{z}$ than $t_{x/y}$ because the electron diffuses slower along the transport direction, i.e., $D_{z}<D_{x/y}$. In the limit of $t_{z}/t_{x/y}\rightarrow\infty$, the electrons have fully traversed the $L_{x/y}$ several times before leaving the central region. Thus, the transmission after ergodic scattering shows only minimal conductance fluctuations.
@@
We conclude this section by noting the equivalent roles of the spatial size $L_{\alpha}$ and the anisotropic parameters $M_{\alpha}\propto D_{\alpha}$ in Eq. 10. Another way to understand the equivalence is to absorb the anisotropy parameters in the Hamiltonian in Eq. 2 into the Hamiltonian in Eq. 3 by rescaling the lattice constant in the hopping term in Eq. 5 as $\hat{a}_{\alpha}\equiv a_{\alpha}/\sqrt{M_{\alpha}}$. This way, one can interpret that the system has isotropic dispersion but with anisotropic hopping terms due to the choice of lattice constants. Meanwhile, the system size under this interpretation is rescaled to $\hat{L}_{\alpha}\equiv L_{\alpha}/\sqrt{M_{\alpha}}$, that is, the same number of lattice sites multiplied by the lattice constant $\hat{a}_{\alpha}$. The fact that band anisotropy and spatial anisotropy are transferable justifies their equivalence. In analogy with the dependence of $c_{d}$ on spatial anisotropy, we define the regions of sensitivity ($M_{z}/M_{y}\geq 1$) and insensitivity($M_{z}/M_{y}<1$) for $\tilde{c}_{d}$ in Fig. 1, as separated by the dashed line. Within the region of (in)sensitivity, $\tilde{c}_{d}$ is (in)sensitive to the changes in anisotropy parameter $M_{z}/M_{y}$. In summary, in the presence of both anisotropies in $L_{\alpha}$ and $M_{\alpha}$, it is convenient to use the relative diffusion time $t_{z}/t_{x/y}$ that converts all these anisotropies to estimate the UCF amplitude in the diffusive free-electron gas.